OpenStudy (anonymous):

Given the following information, determine the Keq for the following reaction: COCl2 (g) 2-way arrow CO (g) + Cl2 (g) Substance Concentration (M) equilibrium CoCl2 1.8 x 10-3 CO 1.2 x 10-2 Cl2 1.2 x 10-2 a. 1.0 x 103 b. 12.5 c. 8.0 x 10-2 d. 9.6 x 10-4

OpenStudy (anonymous):

From what I remember... \(\sf \large K_{eq}=\frac{[CO][Cl_2]}{[COCl_2]}\) just plug in the given values and that would be the equilibrium constant

OpenStudy (anonymous):

\(\sf \large K_{eq}=\frac{[CO][Cl_2]}{[COCl_2]}\\\sf \large K_{eq}=\frac{[1.2 x 10^{-2}][1.2 x 10^{-2}]}{[1.8 x 10^{-3}]}\\K_{eq}=?\)